Sequential (Step) Method of Support Department Cost Allocation

Valron Company has two support departments, Human Resources and General Factory, and two producing departments, Fabricating and Assembly.


Support Departments
Producing Departments
Human
Resources
General
Factory
Fabricating Assembly
Direct costs $160,000 $340,000 $114,600 $93,000
Normal activity:
Number of employees 60 80 170
Square footage 1,000 5,700 13,300

The costs of the Human Resources Department are allocated on the basis of number of employees, and the costs of General Factory are allocated on the basis of square footage. Now assume that Valron Company uses the sequential method to allocate support department costs. The support departments are ranked in order of highest cost to lowest cost.

Required:

1. Calculate the allocation ratios (rounded to four significant digits) for the four departments using the sequential method. If an amount is zero, enter "0". Use the rounded values for subsequent calculations.

Proportion of Driver Used by
Human Resources General Factory Fabricating Assembly
Human Resources
General Factory

2. Using the sequential method, allocate the costs of the Human Resources and General Factory departments to the Fabricating and Assembly departments. If an amount is zero, enter"0".

Support Departments Producing Departments
Human Resources General Factory Fabricating Assembly
Direct costs $ $ $ $
Allocate:
  General Factory            
  Human Resources            
Total after allocation $ $ $
Answer #1

Answers

  • [1]

HR

GF

Fabricating

Assembly

Human Resources [80+170 = 250]

0.00000

0.00000

0.3200

0.6800

General factory [1000 + 5700 + 13300 = 20000]

0.0500

0.0000

0.2850

0.6650

--working

HR

GF

Fabricating

Assembly

Human Resources [80+170 = 250]

0

0

=80/250

=170/250

General factory [1000 + 5700 + 13300 = 20000]

=1000/20000

0

=5700/20000

=13300/20000

  • [2]

HR

GF

Fabricating

Assembly

Direct Overhead cost

$160,000

$340,000

$114,600

$93,000

Allocate:

General factory

$17,000

($340,000)

$96,900

$226,100

Human Resource

($177,000)

$0

$56,640

$120,360

Total after allocation

$0

$0

$268,140

$439,460

--Working

HR

GF

Fabricating

Assembly

Direct Overhead cost

160000

340000

114600

93000

Allocate:

General factory

=340000*0.05

-340000

=340000*0.285

=340000*0.665

Human Resource

=-160000-17000

0

=177000*0.32

=177000*0.68